# Privacy Preserving Mechanisms for Coordinating Airspace Usage in Advanced Air Mobility

### v1

@@ 2 Model of Advanced Air Mobility
Consider the problem of scheduling Advanced Air Mobility (AAM) vehicles, such as drones and eVTOLs, while facilitating novel AAM services in urban airspace. In our model, we segment the urban airspace into contiguous regions or sectors. We denote the set of regions as $\mathcal{R}$. The spatial layout of the airspace is represented as a graph $\mathcal{G}=(\mathcal{R},\mathcal{E})$, where $\mathcal{R}$ corresponds to the set of vertices of graph $\mathcal{G}$, and $\mathcal{E}\subseteq\mathcal{R}\times\mathcal{R}$ represents the set of edges that connect contiguous regions, indicating the feasible movement of AAM vehicles. We model AAM vehicles arriving dynamically throughout the day, dividing the entire day into $T$ time intervals (for example, 1-minute intervals). The set of AAM vehicle requests arriving in the system [^1] We assume that $T$ is sufficiently large to allow us to round the arrival time of each AAM vehicle to the beginning of the corresponding time interval. at time $t$ is denoted by $U(t)$. Every AAM vehicle $u\in U(t)$ has a desired “menu” of time-trajectories, also known as “air-corridors,” one of which it prefers to follow in the airspace. We denote the requested menu for the AAM vehicle $u\in U(t)$ as $M_{u}$ [^2] The set $M_{u}$ includes an option for the vehicle to drop out of the system, denoted by $\varnothing$. (to be defined formally later). Due to capacity constraints, it may not be feasible to allocate the most preferred time-trajectory to all AAM vehicles, as doing so might violate airspace constraints. Specifically, each region has a limit on the number of vehicles that can simultaneously arrive, depart, and stay in that region at any given time  [^3] This constraint arises from safety concerns that limit the number of vehicles that can be autonomously de-conflicted in a confined space (Bauranov and Rakas, 2021).. We denote $C^{\textsf{arr}}(r,t),C^{\textsf{dep}}(r,t)$ and $C^{\textsf{stay}}(r,t)$ to be the number of vehicles that can arrive, depart and stay in region $r\in\mathcal{R}$ at any instance of time $t$, respectively [^4] The arrival and departure constraints are primarily required for airspace regions comprising of vertiports/launch-pads..
A service provider (SP) is responsible for managing the airspace and allocating each AAM vehicle a feasible time-trajectory from its menu while ensuring compliance with airspace constraints. To facilitate this process, we introduce the framework of a time-extended graph, which is essential for integrating arrival, departure, and transit constraints when allocating time-trajectories to AAM vehicles.
Figure 2: A time-trajectory of a drone delivering a package in an urban setting. The drone starts from the launch pad V1 (Sector 1) and needs to drop a package in Sector 5 before returning. Here, we have shown a simple trajectory that moves between regions in one time step, but in general, such trajectories can remain in any region for multiple rounds.
@@
    • ~$\tilde{\mathcal{E}}^{(1)}:=\cup_{t=1}^{T}\{(\nu^{\textsf{arr}}(r,t),\nu(r,t))\}$. Any edge of the type $(\nu^{\textsf{arr}}(r,t),\nu(r,t))$ has capacity $C^{\textsf{arr}}(r,t)$.
@@
    • ~$\tilde{\mathcal{E}}^{(3)}:=\cup_{t=1}^{T-1}\{(\nu(r,t),\nu(r,t+1))\}$. Any edge of the type $(\nu(r,t),\nu(r,t+1))$ has the capacity of $C^{\textsf{stay}}(r,t)$.
@@
A simple example describing the time-extended graph is provided in Section A.
The menu for any AAM vehicle $u\in U(t)$ is defined as $M_{u}=R_{u}\cup\{\varnothing\}$, where $R_{u}$ is a subset of paths on the time-extended graph, and $\varnothing$ represents the option to drop out of the system if the AAM vehicle cannot be allocated to a feasible path due to high congestion. Fig. 2 is a schematic illustrating a time-trajectory scenario for a package delivery drone. Each AAM vehicle $u\in U(t)$ derives a (private) value $v_{u,s}\in\mathbb{R}_{\geq 0}$ if it is allocated an option $s\in M_{u}$.
In this work, we develop an efficient algorithm for the SP to allocate airspace regions to AAM vehicles on demand. The algorithm ensures that capacity constraints are respected while eliminating the need for the SP to know the private valuations of the AAM vehicles. In the next subsection, we propose a market-based solution to address the problem of allocating airspace regions to the AAM vehicles.
3 Market Mechanisms for Airspace Management
We propose that the service provider (SP) allocates regions to AAM vehicles in a receding horizon manner by periodically gathering demand from UAVs and distributing the available capacity of airspace regions using an auction mechanism (to be defined shortly). In each round of the receding horizon, the SP grants each participating vehicle a certain amount of ‘air credits’, which act as a form of artificial currency for purchasing access to the airspace in that round. Specifically, for any given horizon, the SP assigns each vehicle a budget of $w_{u}\geq 0$ air credits [^5] The variation in budgets is due to two factors: (i) the SP prioritizing disaster relief and emergency service vehicles, and (ii) the savings individual AAM vehicles accumulate from previous rounds of interaction. to each AAM vehicle $u$. The SP charges an anonymous payment of $p_{e}$ for the use of each edge $e\in\tilde{\mathcal{E}}$. Given these constraints, each AAM vehicle $u\in U(t)$ selects an option $s\in M_{u}$ that maximizes its value under the allocated budget.
If an AAM vehicle is not allocated during a round, it can convert its unused air credits into an “outside option”, denoted by o, for future use. The SP assigns a price $p_{\textsf{o}}\geq 0$ per unit of the outside option, while the utility that each vehicle derives from this outside option is represented by $v_{u,\textsf{o}}>0$.
In what follows, we only focus on one round of the auction mechanism, with the understanding that this mechanism is implemented in a receding horizon manner. The set of AAM vehicles demanding access in one round of auctions is denoted by $U$. We define $x_{u,e}\in\{0,1\}$ to denote whether an AAM vehicle $u$ is using edge $e\in\tilde{\mathcal{E}}$ and $x_{u,\varnothing}\in\{0,1\}$ if the AAM vehicle $u$ has dropped out of the system. Finally, we use $x_{u,\textsf{o}}$ to denote the amount of outside options that the AAM vehicle consumes for future use. For concise notation, we define $\mathbf{x}_{u}=(x_{u,e})_{e\in\tilde{\mathcal{E}}}$ and $\bar{\mathbf{x}}_{u}=[\mathbf{x}_{u}^{\top},x_{u,\textsf{o}},x_{u,\varnothing}]^{\top}$. Consequently, the overall utility derived by AAM vehicle $u$ is given by
@@ 3 Market Mechanisms for Airspace Management
where $e^{\ast}(s)$ denotes the departing edge [^6] In our framework we use the convention that the AAM vehicles place all their valuation of a route on the first edge on the time-extended graph that goes out from the origin node. This is an edge of type ~$\tilde{\mathcal{E}}^{(4)}$ in Definition 2.1. on the route $s$. Upon observing the price $\mathbf{p}=(p_{e})_{e\in\tilde{\mathcal{E}}}$, each AAM vehicle solves the following optimization problem:
@@
where $\tilde{\mathbf{a}}_{u}\in\mathbb{R}^{|\tilde{\mathcal{E}}|}$ is such that the constraint (2b) enforces $\sum_{s\in R_{u}}x_{u,e^{\ast}(s)}+x_{u,\varnothing}=1$, indicating that the AAM vehicle $u$ will either select a path in $R_{u}$ or will drop out. The matrix $\tilde{\mathbf{A}}_{u,s}\in\mathbb{R}^{K\times(|\tilde{\mathcal{E}}|)}$ in (2c) represents two types of constraints: (i) $\textbf{A}_{u}\mathbf{x}_{u}=\mathbf{0}$, where $\textbf{A}_{u}\in\mathbb{R}^{|\tilde{\mathcal{R}}|\times|\tilde{\mathcal{E}}|}$ is an incidence matrix of the time-extended graph encoding flow-balance constraints at each node in $\tilde{\mathcal{G}}$; and (ii) $\mathbf{B}_{u,s}\mathbf{x}_{u}=0$ for each $s\in R_{u},$ where $\mathbf{B}_{u,s}\in\mathbb{R}^{(K-|\tilde{\mathcal{R}}|)\times|\tilde{\mathcal{E}}|}$ encodes that the flow on any edge connecting two different regions along path $s$ matches the flow on the departing edge $e^{\ast}(s)$. Intuitively, (ii) ensures that any feasible solution that satisfies (2b) and (i) results in a unique edge flow. We present a simple example in Section A to describe these constraints.
@@
The goal of the SP is to set the prices such that the resulting allocation is efficient in the following manner: the prices and consumption levels are at a competitive equilibrium.
@@
    For every $u\in U,$ $\bar{\mathbf{x}}_{u}^{\ast}$ is an optimal solution of (2) with price set to $\mathbf{p}^{\ast}$;
@@
In general, a competitive equilibrium may not always exist (Budish, 2011). Therefore, we introduce a relaxed version of competitive equilibrium, where we relax the requirement that allocations are integral.
@@
$(\bar{\mathbf{x}}^{\ast},\mathbf{p}^{\ast})$ is called the fractional-competitive equilibrium if all conditions in Definition 3.1 are satisfied, except that in Definition 3.1-(i) the integral constraint (2d) is relaxed to a positivity constraint.
This relaxation is inspired by the competitive equilibrium framework studied in the literature on Fisher markets with linear constraints (Jalota et al., 2023). In the following, we demonstrate that a fractional competitive equilibrium always exists and can be computed as the solution to a fixed-point problem. Finally, in the next section, we leverage this property to develop a two-step algorithmic procedure that produces an integral allocation and prices approximating a competitive equilibrium, without requiring knowledge of the private valuations of the AAM vehicles.
Existence and Computation of Fractional Competitive Equilibrium
Next, we establish the following existence result:
@@
The proof builds on the result establishing the existence of a competitive equilibrium in Fisher markets with auxiliary inequality constraints (Jalota et al., 2023). Specifically, our proof accounts for auxiliary equality constraints that arise from (2b)-(2c). A detailed proof of Proposition 3.3 is provided in subsection B.1. Next, we present an optimization-based framework to compute a fractional-competitive equilibrium. Consider the following optimization problem, parametrized by $\omega\in\mathbb{R}_{\geq 0}^{|U|}:$
@@
$\quad\tilde{\mathbf{a}}_{u}^{\top}\mathbf{x}_{u}+x_{u,\varnothing}=1$ (3c)
$\quad\tilde{\mathbf{A}}_{u}\mathbf{x}_{u}=\mathbf{0}$ (3d)
@@
where the first term in (3a) represents the “budget-adjusted” weighted geometric mean of the utilities of all AAM vehicles, while the second term accounts for the total expenditure on the outside option. Constraint (3b) enforces the capacity limit on every edge, and constraints (3c)-(3d) are analogous to (2b)-(2c). Additionally, (3e) is a relaxation of (2d). The objective in (3) is related to the ”budget-adjusted social optimization problem” studied in (Jalota et al., 2023) for Fisher markets, with the key difference being the second term, which ensures that a smaller amount of credits is spent on the outside option.
@@
The constraints (3b)-(3e) always have a feasible solution. Furthermore, for any $\mathbf{\omega}\geq\mathbf{0}$, (3) is a convex optimization problem.
The proof of this result follows from the fact that the constraint set in (3) is a polytope. Moreover, the objective is a concave function as $f_{u}(\cdot)$ is a linear function with positive coefficients.
For any $\omega\in\mathbb{R}^{|U|}_{\geq 0}$, let $\mathbf{p}^{\dagger}(\omega)$ denote an optimal dual multiplier corresponding to the constraint (3b), $\lambda^{\dagger}(\omega)$ denote an optimal dual multiplier corresponding to the constraint (3c), and $\bar{\mathbf{x}}^{\dagger}(\omega)$ denote an optimal solution to (3).
@@
Suppose there exists $\omega^{\ast}\in\mathbb{R}^{|U|}_{\geq 0}$ which is a fixed point of the mapping $\omega\mapsto\lambda^{\dagger}(\omega)$. Then $(\bar{\mathbf{x}}^{\dagger}(\omega^{\ast}),\mathbf{p}^{\dagger}(\omega^{\ast}))$ is a fractional-competitive equilibrium.
The proof relies on the convexity of the optimization problems (3) (Lemma 3.4) and (2) (after relaxing the constraint (2d) to be fractional), along with matching the KKT conditions for optimality. The detailed proof of Proposition 3.5 is provided in Section B.2.
@@
The proof of Proposition 3.5 can be extended to settings where the fixed point may not exist. Suppose there exists $\omega^{\ast}\in\mathbb{R}^{|U|}_{\geq 0}$ such that, for each $u\in U$, $\omega^{\ast}_{u}-\lambda^{\dagger}_{u}(\omega^{\ast})=\epsilon_{u}$ for some $\epsilon_{u}\in\mathbb{R}$ ensuring that $w_{u}+\epsilon_{u}\geq 0$. Then $(\bar{\mathbf{x}}^{\dagger}(\omega^{\ast}),\mathbf{p}^{\dagger}(\omega^{\ast}))$ is a fractional-competitive equilibrium of a market where, for each $u\in U$, the budget is adjusted to $w_{u}+\epsilon_{u}$.
4 Algorithm
In this section, we outline our algorithmic procedure for the service provider (SP) to allocate airspace regions efficiently to AAM vehicles, without any knowledge of their private valuations. Our approach unfolds in two stages. First, we introduce an algorithm that solves the fixed-point equation from Proposition 3.5 to compute the fractional-competitive equilibrium in a distributed manner. Next, using the prices and fractional allocations derived from this step, the SP generates a ranked list of AAM vehicles. This ranking allows the SP to achieve an integral allocation by successively assigning regions to AAM vehicles according to the ranking. In the following subsections, we elaborate on each of these steps.

### v2

@@ 2 Model of Advanced Air Mobility
Consider the problem of allocating airspace to Advanced Air Mobility (AAM) vehicles, such as drones and eVTOLs, which enable novel AAM services in urban environments. In our model, we segment the urban airspace into contiguous regions or sectors, denoted by $\mathcal{R}$. The spatial configuration of the airspace is represented as a graph $\mathcal{G}=(\mathcal{R},\mathcal{E})$, where $\mathcal{R}$ corresponds to the set of vertices, and $\mathcal{E}\subseteq\mathcal{R}\times\mathcal{R}$ represents the set of edges that connect adjacent regions, indicating feasible movements for AAM vehicles between regions. To account for the temporal dimension, we divide the entire day into $T$ time-steps (with each time-step comprising of $\tau$ seconds of time). AAM vehicles arrive dynamically, each requesting access to the airspace.
Each AAM vehicle has a feasible set of time-trajectories (also referred to as “air corridors”) that it can utilize within the airspace. Each vehicle can independently determine the set of feasible trajectories by accounting for energy consumption, travel time, and other operational factors. Furthermore, the feasible set only includes time-trajectories that start before the vehicle’s takeoff; mid-flight trajectory changes are not permitted. Fig. 2 illustrates a schematic representation of a time-trajectory for a package delivery scenario. For every feasible trajectory, the AAM vehicle generates a private valuation that may differ across trajectories, reflecting its preferences over trajectories.
In our model, we assume that each region has a limit on the number of vehicles that can simultaneously arrive, depart, or remain in that region at any given time [^1] This constraint arises from safety concerns that limit the number of vehicles that can be autonomously de-conflicted in a confined space (Bauranov and Rakas, 2021).. Due to capacity constraints, it may not be feasible to allocate each AAM vehicle its most preferred time-trajectory, as this could violate airspace constraints. In such a scenario, a service provider (SP) is typically responsible for managing the airspace and assigning each AAM vehicle a feasible time-trajectory while ensuring compliance with airspace constraints. To facilitate this process, we introduce the framework of a time-extended graph, which is essential for integrating arrival, departure, and transit constraints when allocating time-trajectories to AAM vehicles.
Figure 2: A time-trajectory of a drone delivering a package in an urban setting. The drone starts from the launch pad V1 (Sector 1) and needs to drop a package in Sector 5 before returning. Here, we have shown a simple trajectory that moves between regions in one time-step, but in general, such trajectories can remain in any region for multiple time-steps.
@@
    • ~$\tilde{\mathcal{E}}^{(1)}:=\cup_{t=1}^{T}\{(\nu^{\textsf{arr}}(r,t),\nu(r,t))\}$. Any edge of the type $(\nu^{\textsf{arr}}(r,t),\nu(r,t))$ has capacity [^2] The arrival and departure constraints are primarily required for airspace regions comprising of vertiports/launch-pads. $C^{\textsf{arr}}(r,t)$.
@@
    • ~$\tilde{\mathcal{E}}^{(3)}:=\cup_{t=1}^{T-1}\{(\nu(r,t),\nu(r,t+1))\}$. Any edge of the type $(\nu(r,t),\nu(r,t+1))$ has capacity $C^{\textsf{stay}}(r,t)$.
@@
Any time-trajectory of an AAM vehicle is a path on this time-extended graph. A simple example describing the time-extended graph and time-trajectories is provided in Appendix A.
In this work, we propose a market-based mechanism for the service provider (SP) to allocate capacity-constrained airspace infrastructure to dynamically arriving AAM vehicles. Our mechanism ensures that: $(i)$ capacity constraints of the airspace are strictly satisfied, $(ii)$ each AAM vehicle receives an (approximately) optimal allocation according to its preferences, and $(iii)$ AAM vehicles are not required to disclose their private valuations to the SP. The detailed design and implementation of this mechanism are discussed in the following section.
3 High-level Overview of Market-based Mechanism
We propose an auction-based approach that allocates airspace to dynamically arriving AAM vehicles in a receding-horizon fashion by periodically allocating the available airspace capacity through an auction mechanism. In particular, over a total duration of $T$ time-steps (with each time-step comprising of $\tau$ seconds of time), the service provider (SP) conducts $I$ auctions. In every auction, the SP allocates the airspace to AAM vehicles participating in that auction.
Figure 3: A schematic depiction of the receding horizon approach.
The allocation of the $i$-th auction is determined at time-step $t_{i}=(i-1)\lfloor T/I\rfloor+1.$ The set of vehicles participating in the $i$-th auction comprises of new vehicles that have requested airspace access to SP after the $(i-1)$-th auction (i.e. after time-step $t_{i-1}$) but also those that were not allocated in previous auctions. Each AAM vehicle has a feasible set of time-trajectories that it can follow [^3] Note that the time-trajectory allocated in auction $i$ may start at time-step $t_{i}$ and can end at time-step $T$.. In each auction, the SP assigns a time-trajectory to each participating AAM vehicle from its feasible set444 Note that the feasible set of trajectories of each vehicle include an option $\varnothing$ which indicates that the vehicle will not take-off in that auction (i.e remain unallocated), which makes sure that our problem always has a feasible allocation.. It is important to emphasize that the flight trajectory of each AAM vehicle is finalized before takeoff, and they are not allowed to participate in subsequent auctions to modify their trajectory mid-flight. Finally, the SP updates the remaining airspace capacity before initiating the next auction. A schematic of the receding-horizon approach is provided in Fig. 3.
At a high level, in each auction, the SP allocates a certain amount of “air credits” to each participating AAM vehicle. These air credits, along with other credits they have from past auctions, act as an artificial currency for purchasing airspace access during that auction. Additionally, the SP imposes a payment (in air credits) for the use of each edge in the time-extended graph. AAM vehicle can utilize any leftover budget in future auctions to purchase access to the airspace.
The main design component of this auction mechanism is to determine these prices in a way that allows each AAM vehicle to afford an (approximately) optimal time-trajectory—one that maximizes its utility within its allocated budget—while ensuring that the overall airspace allocation adheres to capacity constraints. Moreover, these prices must be computed in a manner that preserves the privacy of AAM vehicles, ensuring that they do not need to disclose their private valuations. In Section 4, we formally present one such auction mechanism that is used in the receding horizon approach discussed above, assuming that the SP has access to private valuations. Additionally, we study the theoretical properties of the proposed auction mechanism. In Section 5, we relax the assumption that the SP knows the private valuations of AAM vehicles and develop an algorithmic approach (Algorithm 1-2) to implement the proposed auction mechanism without requiring this knowledge.
4 Auction Mechanism: Design and Analysis
In this section, we formally describe the elements of our proposed auction mechanism, along with theoretical guarantees, assuming that the service provider (SP) has access to the private valuations of each AAM vehicle.
[…]
We define $x_{u,e}\in\{0,1\}$ to indicate whether an AAM vehicle $u$ is using edge $e\in\tilde{\mathcal{E}}$, and $x_{u,\varnothing}\in\{0,1\}$ to indicate whether the AAM vehicle $u$ has dropped out of the system. Furthermore, each AAM vehicle has the option to convert its unused budget into an “outside option” for future auctions666 The budget converted into the outside option is carried forward and added to the AAM vehicle’s budget in future auctions.. We use $x_{u,\textsf{o}}$ to denote the amount of outside options that the AAM vehicle consumes in any auction. For concise notation, we define $\mathbf{x}_{u}=(x_{u,e})_{e\in\tilde{\mathcal{E}}}$ and $\bar{\mathbf{x}}_{u}=\begin{bmatrix}\mathbf{x}_{u}^{\top},x_{u,\textsf{o}},x_{u,\varnothing}\end{bmatrix}^{\top}.$
The utility derived by any vehicle $u\in U$ from selecting a route $s\in R_{u}$ is denoted by $v_{u,s}\in\mathbb{R}_{+}$, selecting the option $\varnothing$ is denoted by $v_{u,\varnothing}\in\mathbb{R}_{+}$, and per unit consumption of outside option is given by $v_{u,\textsf{o}}\in\mathbb{R}_{+}$. Therefore, the overall utility derived by AAM vehicle $u$ is given by
@@ 4 Auction Mechanism: Design and Analysis
where $e^{\ast}(s)$ denotes the departing edge [^7] In our framework, we use the convention that the AAM vehicles place all their valuation of a route on the first edge on the time-extended graph that goes out from the origin node. This is an edge of type ~$\tilde{\mathcal{E}}^{(4)}$ in Definition 2.1. Additionally, we add a constraint (cf. (3b)) that ensures that if a vehicle selects a departing edge corresponding to a route, then all edges on that route will be selected. on the route $s$.
The SP charges a price $p_{e},$ for any AAM vehicle using the edge $e\in\tilde{\mathcal{E}}$, and a payment $p_{\textsf{o}}\geq 0$ for the consumption per unit of the outside option.
Upon observing the prices, each AAM vehicle solves the following optimization problem:
@@
where $\tilde{\mathbf{a}}_{u}\in\mathbb{R}^{|\tilde{\mathcal{E}}|}$ is such that the constraint (2b) enforces $\sum_{s\in R_{u}}x_{u,e^{\ast}(s)}+x_{u,\varnothing}=1$, indicating that the AAM vehicle $u$ will either select a path in $R_{u}$ or will drop out. The matrix $\tilde{\mathbf{A}}_{u,s}\in\mathbb{R}^{K\times(|\tilde{\mathcal{E}}|)}$ in (2c) represents two types of constraints: (i) $\textbf{A}_{u}\mathbf{x}_{u}=\mathbf{0}$, where $\textbf{A}_{u}\in\mathbb{R}^{|\tilde{\mathcal{R}}|\times|\tilde{\mathcal{E}}|}$ is an incidence matrix of the time-extended graph encoding flow-balance constraints at each node in $\tilde{\mathcal{G}}$; and (ii) $\mathbf{B}_{u,s}\mathbf{x}_{u}=0$ for each $s\in R_{u},$ where $\mathbf{B}_{u,s}\in\mathbb{R}^{(K-|\tilde{\mathcal{R}}|)\times|\tilde{\mathcal{E}}|}$ encodes the constraint that the flow on any edge connecting two different regions along path $s$ matches the flow on the departing edge $e^{\ast}(s)$. Intuitively, (ii) ensures that any feasible solution that satisfies (2b) and (i) results in a unique edge flow. We present a simple example in Appendix A to describe these constraints.
@@
The goal of the SP is to set the prices such that the resulting allocation is a competitive equilibrium:
@@
    For every $u\in U,$ $\bar{\mathbf{x}}_{u}^{\ast}$ is an optimal solution of (2) with prices set to $\mathbf{p}^{\ast}$;
@@
We call $\bar{\mathbf{x}}^{\ast},\mathbf{p}^{\ast}$ to be the market clearing allocation and market clearing prices, respectively. In general, a competitive equilibrium may not always exist (Budish, 2011). Therefore, we introduce a relaxed version of competitive equilibrium, where we relax the requirement that allocations are integral.
@@
$(\bar{\mathbf{x}}^{\ast},\mathbf{p}^{\ast})$ is called a fractional-competitive equilibrium if all conditions in Definition 4.1 are satisfied, except that in Definition 4.1-(i) the integral constraint (2d) is relaxed to a positivity constraint.
This relaxation is inspired by the competitive equilibrium framework studied in the literature on Fisher markets with linear constraints (Jalota et al., 2023).
In what follows, we demonstrate that a fractional competitive equilibrium always exists and can be computed as the solution to a fixed-point problem. Next, in Section 5, we leverage this property to develop a two-step algorithmic procedure that produces an integral allocation and prices approximating a competitive equilibrium, without requiring knowledge of the private valuations of the AAM vehicles.
4.1 Existence and Computation of Fractional Competitive Equilibrium
We state the following existence result about the fractional competitive equilibrium.
@@ 4.1 Existence and Computation of Fractional Competitive Equilibrium
The proof builds on the result establishing the existence of a competitive equilibrium in Fisher markets with auxiliary inequality constraints (Jalota et al., 2023). Specifically, our proof accounts for auxiliary equality constraints that arise from (2b)-(2c). A detailed proof of Proposition 4.3 is provided in Appendix B.1.
Next, we present a computational framework that can be used by the service provider for computing a fractional-competitive equilibrium, if they know the private valuations of all AAM vehicles. Consider the following optimization problem, parametrized by $\omega\in\mathbb{R}_{\geq 0}^{|U|}:$
@@
$\quad\tilde{\mathbf{a}}_{u}^{\top}\mathbf{x}_{u}+x_{u,\varnothing}=1\quad\forall\ u\in U$ (3c)
$\quad\tilde{\mathbf{A}}_{u}\mathbf{x}_{u}=\mathbf{0}\quad\forall\ u\in U$ (3d)
@@
where the first term in (3a) represents the “budget-adjusted” weighted geometric mean of the utilities of all AAM vehicles, while the second term accounts for the total expenditure on the outside option. Constraint (3b) enforces the capacity limit on every edge, and constraints (3c)-(3d) are analogous to (2b)-(2c). Additionally, (3e) is a relaxation of the integrality constraint in (2d). The objective in (3) is related to the “budget-adjusted social optimization problem” studied in (Jalota et al., 2023) for Fisher markets, with the key difference being the second term, which ensures that a smaller amount of credits is spent on the outside option. On an intuitive level, the weighted geometric mean structure of the objective in (3a) can be seen as finding an allocation that balances the trade-offs between different AAM vehicles’ utilities, weighted by their budget adjustments. It ensures that an improvement in a vehicle’s utility contributes to the overall objective in proportion to its market power. At the optimal point, this results in a fair and efficient allocation of airspace among AAM vehicles. On a more technical level, the weighted geometric mean is fundamental in the proof of Proposition 4.5, which ensures that if we can adjust the weights of AAM vehicles in an appropriate manner (through a careful choice of $\omega$) then the optimal dual multiplier of (3b) is the market clearing price and the optimizer of (3) are market clearing allocations.
Before stating Proposition 4.5, we present an important property of (3).
@@
The constraints (3b)-(3e) always have a feasible solution. Furthermore, for any $\omega\in\mathbb{R}^{|U|}_{+}$, (3) is a convex optimization problem.
The proof of this result follows from the fact that the constraint set in (3) is a polytope. Moreover, the objective $f_{u}(\cdot)$ is a linear function with positive coefficients. For any $\omega\in\mathbb{R}^{|U|}_{+}$, let $\mathbf{p}^{\dagger}(\omega)$ denote an optimal dual multiplier corresponding to the constraint (3b), $\lambda^{\dagger}(\omega)$ denote an optimal dual multiplier corresponding to the constraint (3c), and $\bar{\mathbf{x}}^{\dagger}(\omega)$ denote an optimal solution to (3).
@@
Suppose there exists $\omega^{\ast}\in\mathbb{R}^{|U|}_{+}$ that is a fixed point of the mapping $\omega\mapsto\lambda^{\dagger}(\omega)$. Then $(\bar{\mathbf{x}}^{\dagger}(\omega^{\ast}),\mathbf{p}^{\dagger}(\omega^{\ast}))$ is a fractional-competitive equilibrium.
The proof relies on the convexity of the optimization problems (3) (Lemma 4.4) and (2) (after relaxing the integrality constraint in (2d) to fractional in (3e)), along with matching the KKT conditions for optimality. The detailed proof of Proposition 4.5 is provided in Appendix B.2.
@@
The proof of Proposition 4.5 can be extended to settings where a fixed point may not exist. Suppose there exists $\omega^{\ast}\in\mathbb{R}^{|U|}_{+}$ such that, for each $u\in U$, $\omega^{\ast}_{u}-\lambda^{\dagger}_{u}(\omega^{\ast})=\epsilon_{u}$ for some $\epsilon_{u}\in\mathbb{R}$ ensuring that $w_{u}+\epsilon_{u}\geq 0$. Then $(\bar{\mathbf{x}}^{\dagger}(\omega^{\ast}),\mathbf{p}^{\dagger}(\omega^{\ast}))$ is a fractional-competitive equilibrium of a market where, for each $u\in U$, the budget is adjusted to $w_{u}+\epsilon_{u}$.
5 Algorithmic Design of Auction Mechanism without Private Valuations
In this section, we outline our algorithmic procedure for the service provider (SP) to compute (approximate) competitive equilibria, using Proposition 4.5, without knowing the private valuations of AAM vehicles. Our approach unfolds in two stages. First, we introduce an algorithm that solves the fixed-point equation from Proposition 4.5 to compute the fractional-competitive equilibrium in a distributed manner (cf. Algorithm 1). The SP then generates a ranked list of AAM vehicles using the prices and fractional allocations derived from this step. This ranking allows the SP to achieve an integral allocation by successively assigning regions to AAM vehicles according to the ranking (cf. Algorithm 2). In the following subsections, we elaborate on each of these steps.